OpenStudy (anonymous):

If there are 91 slips of paper with odd or even numbers on them, and the probability of picking a slip with an odd number is 2/7, how many slips of paper have numbers divisible by 2 on them?

myininaya (myininaya):

i would do \[x+y=91, \frac{x}{x+y}=\frac{2}{7}\]

OpenStudy (anonymous):

Well, if the probability of picking an odd number is 2/7, then the probability of picking an even numbers is: \[1-\frac{2}{7} = \frac{5}{7}\] Then to figure out how many of those 91 papers are even we multiply: \[91*\frac{5}{7}\] = 65? (someone check my math)

OpenStudy (anonymous):

Or myininaya's way, either or :P

OpenStudy (anonymous):

your math is correct

myininaya (myininaya):

26 odd 91-26 even

myininaya (myininaya):

gj joe

OpenStudy (anonymous):

thank you :) gj to you as well!